OpenStudy (anonymous):

Change to an improper fraction 9 1/3 9 2/5 6 3/4 6 2/7

OpenStudy (anonymous):

To change to an improper fraction, multiply the whole number by the demoniator of the fraction, then add it to the numerator. For example, for the first one: \[9 \frac{ 1 }{ 3}\] Multiply 9 x 3 to get 27, then add it to the numerator: 27+1=28 \[ \frac{ 28 }{ 3}\] Use the same method for the rest of the problems.